Perform the required operations on the following functions.
Given: f(x) = 3 - x; g(x) = -2x
Find g[f(x)].
A.) 2x - 6
B.) -x - 6
C.) 2x + 3
Please include explanation, thanks. :)

Answer by Cromlix(4381)

Hi there,
f(x) = 3 - x g(x) = -2x
g(f(x)) = -2(3 - x)
g(f(x)) = -6 + 2x
(A) is the answer
Hope this helps :-)